# Published in the proceedings of the 39th Conference on Neural Information Processing Systems (NeurIPS 2025) Workshop: Scaling Environments for Agents (SEA). Additionally accepted for presentation in NeurIPS 2025 Workshop: Embodied World Models for Decision Making

### v2

@@ A.2.2 Scaling the predator-prey environment
Listing 8: Pseudocode for 2D predator-prey task
@@
Listing 9: Pseudocode for predator-prey with an adversary
@@
Listing 10: Pseudocode for multi-organoid predator-prey
A.2.3 Scaling the dynamic interception (Pong) environment
The following pseudocode shows how the principles of the Pong task can be generalized to other classic games like Breakout or adapted for multi-agent competition.
@@ A.2.3 Scaling the dynamic interception (Pong) environment
Listing 11: Pseudocode for breakout task
@@
Listing 12: Pseudocode for two-organoid competitive pong
A.3 LLM prompting examples
This appendix provides detailed examples of the two primary modes for using an LLM to generate experimental protocols: structured JSON output for parameter tuning and full Python script generation for implementing novel logic and curriculum design.
A.3.1 Example 1: JSON-based parameter optimization
@@
A.3.2 Example 2: Full script generation for curriculum design
This approach grants the LLM maximum flexibility to design novel experimental logic, suitable for tasks like curriculum design where the protocol itself must change based on agent performance. By providing the LLM with a detailed API specification, it can compose complex sequences of operations. The prompt in Listing 15 asks the LLM to write a complete Python script for a block of trials for a hypothetical platform. This includes defining the logic for state encoding, action decoding, feedback delivery, and a terminal condition with a suggestion for the next curriculum step. This example uses FinalSpark’s Neuroplatform [Jordan et al., 2024].

### v3

@@ 4.2.2 Environmental scaling and multi-agent systems
    Listing 6: Pseudocode for 2D predator-prey task
  • Introduction of an Adversary: A second mobile entity, an adversary, can be introduced, creating a multi-objective task: approach prey while avoiding the adversary. The adversary’s location would be encoded on a separate sensory channel (e.g., via a distinct stimulation frequency). Its behavior could range from a random walk to actively tracking the agent, requiring the agent’s world model to account for multiple dynamic entities with different behavioral characteristics. See Listing 7.
@@
    Listing 7: Pseudocode for predator-prey with an adversary
  • Multi-Organoid Systems: The framework can be extended to create a simple competitive ecosystem. One organoid could control the prey and a second could control the predator, allowing their policies to co-evolve. This can be further extended to a linear food chain (Organoid A is prey to B, B is prey to C) or a cyclical model, in providing a novel platform to study the emergence of complex, hierarchical strategies in a multi-agent biological system. A pseudocode implementation for a two-organoid system is available in Listing 8.
@@
    Listing 8: Pseudocode for multi-organoid predator-prey
4.3 Environment 3: Dynamic interception task (Pong)
Figure 3: The Pong environment. The organoid receives complex, multi-modal sensory input about the ball’s position (rate-coded distance, spatially-coded location) and must learn a predictive world model of the ball’s trajectory to control its paddle and achieve a successful interception.
4.3.1 Protocol description
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Initialize game state and MEA interface
game = PongGame()
Main game loop
while game.is_active:
# 1. Update game state
game.update_ball_position()
ball_x, ball_y = game.get_ball_position()
# 2. State representation
encode_y_pos_rate_coded(groups_C_D, ball_y)
encode_x_pos_spatial(groups_C_D, ball_x)
# 3. Action decoding
spikes_A = record_spikes(group_A, window=10_ms)
spikes_B = record_spikes(group_B, window=10_ms)
if spikes_A > spikes_B:
game.move_paddle_up()
elif spikes_B > spikes_A:
game.move_paddle_down()
# 4. Feedback
if game.check_collision():
deliver_reward_predictable(groups_C_D)
game.ball_bounce()
elif game.check_miss():
deliver_punishment_unpredictable(groups_C_D)
game.reset_ball()
Listing 9: Pseudocode for pong task
4.3.2 Environmental scaling and generalization
  • Generalization to Other Games: The core paradigm of Pong can be adapted to other classic arcade games to probe different cognitive functions and world modeling capabilities. For instance, Breakout would require the agent’s world model to incorporate the same ball and paddle, as well as a static array of targets, demanding precise spatial targeting to aim the ball. Space Invaders would introduce the complexities of dodging projectiles while timing offensive actions, requiring the world model to track multiple independent agents. A pseudocode implementation for Breakout is provided in Listing 10.
@@ 4.3.2 Environmental scaling and generalization
    Listing 10: Pseudocode for breakout task
  • Competitive Multi-Organoid Systems: Pong provides an ideal testbed for multi-agent competition. Two distinct organoid-MEA systems can be linked to the same game instance, each controlling one of the paddles. This setup facilitates the direct study of competitive co-adaptation, where each biological agent must build a world model that includes the policy of its opponent. The evolution of rally lengths, shot patterns, and win/loss rates would serve as powerful metrics for inter-agent learning and strategic model formation (see Listing 11).
@@
Listing 11: Pseudocode for two-organoid competitive pong
To overcome the limitations of manual experimental design and to systematically explore the vast parameter space of environments and training protocols, we propose an automated framework where an LLM functions as a meta-controller or AI Dungeon Master. This approach directly aligns with the theme of leveraging generative models and LLMs, applying the LLM as a tool to orchestrate the generation of the agent’s training world and curriculum. This creates a high-throughput system for discovering optimal training strategies and environmental configurations that might be non-obvious to human researchers. A pseudocode implementation of this meta-learning loop is provided in Listing 12.
4.4.1 Framework description
  1. 1.
  2. 2.
    Protocol Generation: The LLM processes the prompt and generates a novel experimental protocol. This can be done in two primary modes: generating a structured JSON object with specific experimental variables to be plugged into a template script, or generating an entire Python script from scratch that defines novel experimental logic and curriculum stages. This generative capacity allows the LLM to move beyond simple parameter tuning to design entirely new tasks. Detailed examples are provided in Section 4.4.2.
  3. 3.
  4. 4.
  5. 5.
Initialize dataset and load base prompt template
dataset = initialize_database()
prompt_template = load_base_prompt("config/llm_prompt.txt")
Main automation loop
for iteration in range(NUM_ITERATIONS):
# 1. Formulate prompt with historical data
current_organoid_state = get_organoid_baseline()
# 2. LLM generates a new experimental script
generated_script = call_llm_api(prompt)
# 3. Validate the generated script for safety and syntax
# 4. Execute or log error
if is_valid:
# Run the experiment on the MEA platform
[…]
# Store the script and its outcomes in the dataset
else:
# Log the invalid script and the reason for failure
# 5. Periodically refine the LLM’s instructions
if (iteration + 1) % BATCH_SIZE == 0:
prompt_template = refine_prompt_based_on_results(dataset)
# Optional: Trigger fine-tuning if sufficient data is collected
# if len(dataset) > FINETUNE_THRESHOLD:
# launch_finetuning_job(dataset)
Listing 12: Pseudocode for LLM-driven experiment automation
4.4.2 LLM prompting examples
This section provides detailed examples of the two primary modes for using an LLM to generate experimental protocols: structured JSON output for parameter tuning and full Python script generation for implementing novel logic and curriculum design.
Example 1: JSON-based parameter optimization
@@ 4.4.2 LLM prompting examples
Example 2: Full script generation for curriculum design
This approach grants the LLM maximum flexibility to design novel experimental logic, suitable for tasks like curriculum design where the protocol itself must change based on agent performance. By providing the LLM with a detailed API specification, it can compose complex sequences of operations. The prompt in Listing 15 asks the LLM to write a complete Python script for a block of trials for a hypothetical platform. This includes defining the logic for state encoding, action decoding, feedback delivery, and a terminal condition with a suggestion for the next curriculum step. This example uses FinalSpark’s Neuroplatform (Jordan et al., 2024).
@@
5 Discussion
5.1 Broader implications and future work
5.2 Limitations and challenges
Despite its significant promise, this approach has its challenges. The biological substrate introduces inherent variability, limited lifespans, and developmental stochasticity, making reproducibility a central concern that must be addressed with robust experimental design and statistical analysis. This variability arises from multiple sources, including genetic and epigenetic differences between iPSC lines (Mancinelli et al., 2025) and the stochastic nature of self-organization, particularly in unguided protocols (Lancaster and Knoblich, 2014). The limited lifespans and long-term health of large organoids are also a practical constraint, as the tissue’s non-vascularized core is susceptible to hypoxia and necrosis, which can introduce experimental artifacts (Zhang et al., 2022). The organoid itself remains a profound black box; while we can measure network-level plasticity, deciphering the specific computational algorithms it implements to learn and represent its world model is a formidable open problem. Furthermore, the information bandwidth of current MEA technology, while improving, is still orders of magnitude lower than that of biological nervous systems, potentially constraining the complexity of the world models that can be learned.
6 Conclusion
Acknowledgments and Disclosure of Funding
We thank FinalSpark for suggesting the investigation of LLM based automated research.